# Efficient Generation of Multi-partite Entanglement between Non-local Superconducting Qubits using Classical Feedback

### v2

@@ Appendix D Look-up Tables
In this work, look-up tables (LUTs) were utilized for decision branching based on the results of the MCMs. While LUTs are not scalable as system size increases, and will eventually need to be replaced by real-time decoders, they are still viable for small system sizes. The LUTs are constructed by nesting branch statements on QubiC using the distributed processor architecture Fruitwala et al. (2024). These branching statements poll the results of the measured ancillae qubits in a MCM, and then applies a conditional feed-forward operation on data qubit(s) depending on the results of the measurements.
For example, consider the task of preparing an -qubit GHZ state using the protocol introduced in Sec. II. In the stabilizer formalism, a GHZ state is simultaneous eigenstate of and () with eigenvalue . Here, can be used to check the parity of any two neighboring qubits. Thus, the strategy is to initialize the data qubits in the state so that it is an eigenstate of with an eigenvalue of , and then measure the stabilizers in parallel using a MCM on all of the ancillae qubits. The outcome of the stabilizer measurements are then analyzed holistically to determine on which data qubits we should perform a conditional bit-flip to reconstruct the -qubit GHZ state. For example, suppose the data qubits are in the state prior to measurement. By entangling the data qubits pairwise with ancillae qubits, as shown in Fig. 1, we can measure the stabilizers . In the absence of any errors, a measurement of the stabilizers will return the values , where a measurement of () corresponds to a parity of (); this tells us that we should apply a bit-flip ( gate) on the 4th and 5th qubit to recover a GHZ state. Building up a LUT of all possible outcomes allows us to quickly apply conditional bit-flips in real-time, without the latency of having to decode the results on the FPGA or on some other (potentially non-local) classical processor. Though, it is worth noting that a real-time decoder of parity measurements could be implemented by analyzing the parity measurements sequentially (per clock cycle) and considering each 1 measurement as a domain wall which takes us from apply nothing to apply , and back again.
@@ Appendix E Measurement-Induced Dephasing
Mid-circuit measurements (MCMs) can lead to unwanted dephasing on nearby idling spectator qubits. This can happen in two different ways: firstly, when an ancilla qubits is measured, the static coupling Mitchell et al. (2021) between coupled qubits can dephase the spectator qubit during the measurement process Jurcevic and Govia (2022) if the ancilla qubit undergoes a measurement-induced state transition Dumas et al. (2024). This only results in dephasing between qubits that are directly coupled, and assumes that the coupling cannot be turned off during the MCM. This is the case for the qubits used in this experiment, as they are coupled via fix-frequency resonators in a ring geometry (see Fig. 5a). However in systems with tunable couplers, this coupling can be largely suppressed, which would help alleviate the dephasing seen by the spectator qubit.
@@
The other way in which MCMs can dephase spectator qubits is via unintentional direct measurement, i.e., measurement-induced dephasing Gambetta et al. (2006). When the readout resonator frequencies of nearby qubits are not spectrally well-resolved, the measurement of one qubit can inadvertently result in the measurement of other qubits. This is the case for multiple pairs of qubits on the current device (see Fig. 5c). For example, the readout resonators coupled to Q0 and Q1 are within 4 MHz of each other, and the readout resonators coupled to Q3, Q4, and Q5 are within 11 MHz of each other. This is not done intentionally; rather, it is the result of inaccuracies in the fabrication process. This directly limits which qubits can be used as ancillas for MCMs and which can be used as data qubits, because the measurement of the ancilla qubit can result in a complete loss of phase coherence on data qubits.
@@
To measure the impact of MCMs on neighboring spectator qubits, we utilize cycle benchmarking (CB) with a slight variation (see Fig. 7). Here, the interleaved cycle is a MCM on one qubit, with all of the other qubits left idling. However, because the goal is to quantify the impact of MCMs on spectator qubits, we only twirl the spectator qubits, we do not twirl the qubit being measured. A similar strategy was done in Ref. 75 using RB, but RB cannot distinguish the impact of errors along different measurement axes. The utility of CB is that it enables one to quantify how a MCM affects other qubits along , , and . Moreover, we can glean information from both the state-preparation and measurement (SPAM) constant for the exponential decay of each Pauli operator (termed “Pauli decays”), as well as from the exponential fit parameter . If the effect of a MCM on an idle qubit is small, then one should be able to accurately fit and . However, if the effect of a MCM is large enough such that no useable information can be gleaned for a given , then will be small (close to zero) and cannot be trusted (assuming an exponential can even be fit to the data, which is not always the case). This is the case if the MCM completely dephases the idle qubit, for which and (see, e.g., the bottom row of Fig. 9). When an exponential can be fit to the data, gives the probability with which is preserved by the error process. More generally, any Pauli error will attenuate the probability for any non-commuting Pauli operator Hashim et al. (2024). Thus, if a MCM causes a spectator qubit to dephase (which is a -type error), this will affect and (but will not affect ). Therefore, we often observe that , which is a strong indication of measurement-induced dephasing.
@@
Using CB, we systematically characterize the effect of MCMs on all nearest-neighbor idling qubits across our full device (see Figs. 8–11). We compare the effect of MCMs to the case when the spectator qubits idle (with no MCM) for the duration of the measurement, as well as the case when dynamically decoupling (DD; specifically, an XY8 sequence Maudsley (1986); Ahmed et al. (2013)) is applied to the spectator qubits during the MCM. In all cases, we observe that the MCM has at least some effect on the spectator qubits. If the effect is small, we observe that DD helps reduce the dephasing (i.e., increase and ). However, if the effect is large, DD does not help and can even impact the Pauli decay adversely (see, e.g., the bottom row of Fig. 9). This is because DD relies on the assumption that the dephasing noise is quasi-static (i.e., slowly varying) over the course of an experiment, such that one can “echo away” the phase error. However, when a MCM is strong enough to inadvertently measure a nearby qubit, this breaks the underlying assumption of DD, as measurement back-action introduces strong stochastic phase rotations Kockum et al. (2012). In such cases, a DD sequence like XY8 will not return to the qubit to the original state, inevitably introducing - and -type errors (and thus affecting the Pauli- decays).

### v3

@@ Appendix D Look-up Tables
In this work, look-up tables (LUTs) were utilized for decision branching based on the results of the MCMs. While LUTs are not scalable as system size increases, and will eventually need to be replaced by real-time decoders, they are still viable for small system sizes. The LUTs are constructed by nesting branch statements on QubiC using the distributed processor architecture fruitwala2024distributed. These branching statements poll the results of the measured ancillae qubits in a MCM, and then applies a conditional feed-forward operation on data qubit(s) depending on the results of the measurements.
For example, consider the task of preparing an -qubit GHZ state using the protocol introduced in Sec. II. In the stabilizer formalism, a GHZ state is simultaneous eigenstate of and () with eigenvalue . Here, can be used to check the parity of any two neighboring qubits. Thus, the strategy is to initialize the data qubits in the state so that it is an eigenstate of with an eigenvalue of , and then measure the stabilizers in parallel using a MCM on all of the ancillae qubits. The outcome of the stabilizer measurements are then analyzed holistically to determine on which data qubits we should perform a conditional bit-flip to reconstruct the -qubit GHZ state. For example, suppose the data qubits are in the state prior to measurement. By entangling the data qubits pairwise with ancillae qubits, as shown in Fig. 1, we can measure the stabilizers . In the absence of any errors, a measurement of the stabilizers will return the values , where a measurement of () corresponds to a parity of (); this tells us that we should apply a bit-flip ( gate) on the 4th and 5th qubit to recover a GHZ state. Building up a LUT of all possible outcomes allows us to quickly apply conditional bit-flips in real-time, without the latency of having to decode the results on the FPGA or on some other (potentially non-local) classical processor. Though, it is worth noting that a real-time decoder of parity measurements could be implemented by analyzing the parity measurements sequentially (per clock cycle) and considering each 1 measurement as a domain wall which takes us from to apply , and back again.
@@ Appendix E Measurement-Induced Dephasing
Mid-circuit measurements (MCMs) can lead to unwanted dephasing on nearby idling spectator qubits. This can happen in two different ways: firstly, when an ancilla qubits is measured, the static coupling mitchell2021hardware between coupled qubits can dephase the spectator qubit during the measurement process jurcevic2022effective if the ancilla qubit undergoes a measurement-induced state transition dumas2024measurement. This only results in dephasing between qubits that are directly coupled, and assumes that the coupling cannot be turned off during the MCM. This is the case for the qubits used in this experiment, as they are coupled via fix-frequency resonators in a ring geometry (see Fig. 5a). However in systems with tunable couplers, this coupling can be largely suppressed, which would help alleviate the dephasing seen by the spectator qubit.
@@
The other way in which MCMs can dephase spectator qubits is via unintentional direct measurement, i.e., measurement-induced dephasing gambetta2006qubit. When the readout resonator frequencies of nearby qubits are not spectrally well-resolved, the measurement of one qubit can inadvertently result in the measurement of other qubits. This is the case for multiple pairs of qubits on the current device (see Fig. 5c). For example, the readout resonators coupled to Q0 and Q1 are within 4 MHz of each other, and the readout resonators coupled to Q3, Q4, and Q5 are within 11 MHz of each other. This is not done intentionally; rather, it is the result of inaccuracies in the fabrication process. This directly limits which qubits can be used as ancillas for MCMs and which can be used as data qubits, because the measurement of the ancilla qubit can result in a complete loss of phase coherence on data qubits.
@@
To measure the impact of MCMs on neighboring spectator qubits, we utilize cycle benchmarking (CB) with a slight variation (see Fig. 7). Here, the interleaved cycle is a MCM on one qubit, with all of the other qubits left idling. However, because the goal is to quantify the impact of MCMs on spectator qubits, we only twirl the spectator qubits, we do not twirl the qubit being measured. A similar strategy was done in Ref. govia2023randomized using RB, but RB cannot distinguish the impact of errors along different measurement axes. The utility of CB is that it enables one to quantify how a MCM affects other qubits along , , and . Moreover, we can glean information from both the state-preparation and measurement (SPAM) constant for the exponential decay of each Pauli operator (termed “Pauli decays”), as well as from the exponential fit parameter . If the effect of a MCM on an idle qubit is small, then one should be able to accurately fit and . However, if the effect of a MCM is large enough such that no useable information can be gleaned for a given , then will be small (close to zero) and cannot be trusted (assuming an exponential can even be fit to the data, which is not always the case). This is the case if the MCM completely dephases the idle qubit, for which and (see, e.g., the bottom row of Fig. 9). When an exponential can be fit to the data, gives the probability with which is preserved by the error process. More generally, any Pauli error will attenuate the probability for any non-commuting Pauli operator hashim2024practical. Thus, if a MCM causes a spectator qubit to dephase (which is a -type error), this will affect and (but will not affect ). Therefore, we often observe that , which is a strong indication of measurement-induced dephasing.
@@
Using CB, we systematically characterize the effect of MCMs on all nearest-neighbor idling qubits across our full device (see Figs. 8–11). We compare the effect of MCMs to the case when the spectator qubits idle (with no MCM) for the duration of the measurement, as well as the case when dynamically decoupling (DD; specifically, an XY8 sequence maudsley1986modified; ahmed2013robustness) is applied to the spectator qubits during the MCM. In all cases, we observe that the MCM has at least some effect on the spectator qubits. If the effect is small, we observe that DD helps reduce the dephasing (i.e., increase and ). However, if the effect is large, DD does not help and can even impact the Pauli decay adversely (see, e.g., the bottom row of Fig. 9). This is because DD relies on the assumption that the dephasing noise is quasi-static (i.e., slowly varying) over the course of an experiment, such that one can “echo away” the phase error. However, when a MCM is strong enough to inadvertently measure a nearby qubit, this breaks the underlying assumption of DD, as measurement back-action introduces strong stochastic phase rotations kockum2012undoing. In such cases, a DD sequence like XY8 will not return to the qubit to the original state, inevitably introducing - and -type errors (and thus affecting the Pauli- decays).